OpenStudy (anonymous):

Solve -5x-10 <= x + 2 a. x <= 2 b. x <= -2 c. x >= -2 d. x >= 2

OpenStudy (anonymous):

I'm pretty sure it is the answer C, but I need to explain how I got that answer. So I basically need an explanation of how to get the answer.

mathslover (mathslover):

oh k so we have: \[\large{-5x+10 \ge x+2}\] right?

OpenStudy (anonymous):

it is always C

OpenStudy (anonymous):

- 10, not + 10

OpenStudy (anonymous):

1) add \(5x\) 2) subtract \(2\) 3) divide by \(6\)

mathslover (mathslover):

oh k so you can explain like this: \[\large{-5x-10 \le x+2}\] add 10 both sides \[\large{-5x+10-10\le x+2+10}\] \[\large{-5x\le x+12}\] subtract 12 both sides \[\large{-5x-12\le x+12-12}\] \[\large{-5x-12\le x}\] subtract x both sides \[\large{-5x-x-12\le 0}\] \[\large{-6x-12\le 0}\] add 12 both sides \[\large{-6x-12+12\le 0+12}\] \[\large{-6x\le 12}\] divide -6 both sides \[\large{\frac{-6x}{-6}\ge \frac{12}{-6}}\] \[\large{x\ge -2}\]

mathslover (mathslover):

You can explain like this

OpenStudy (anonymous):

i like my method better attack the variables first, constants last same as with linear equations

OpenStudy (anonymous):

So when you divide, the sign switches?

mathslover (mathslover):

right but I put the explanation yes @pathosdebater when we divide a \(negative\) number both sides then the sign changes

OpenStudy (anonymous):

Thanks